# Neutrino mixing parameters and masses from $\Delta(96)\rtimes H_{CP}$ in the tri-direct CP approach

### v1

@@ 3.1 Lepton mixing patterns for $\phi_{\text{atm}},\phi_{\text{sol}}\sim\mathbf{\bar{3}_{0}}$
Moreover, we plot the contour regions for the $3\sigma$ intervals of mixing parameters $\theta_{13}$, $\theta_{23}$, $\delta_{CP}$ and mass ratio $m^{2}_{2}/m^{2}_{3}$ in the plane $r$ and $\eta$ in figure 1. The result for $\theta_{12}$ is not displayed here, because it is related to the reactor angle $\theta_{13}$ by the TM1 mixing sum rule in Eq. (3.12). The JUNO experiment [91] an upcoming medium-baseline reactor neutrino facility, is projected to constrain $\sin^{2}\theta_{12}$ within $[0.3022,0.3118]$ at $3\sigma$ after six years of operation. Complementary tests of $\sin^{2}\theta_{23}$ and $\delta_{CP}$ resolutions (anticipated ranges of $[0.462,0.480]$ and $[1.099,1.255]$ in radians after 15 years, respectively) will be enabled by long-baseline experiments DUNE [92] and T2HK [93]. The predicted neutrino mass sum $\sum_{i=1}^{3}m_{i}$ aligns with cosmological constraints from Planck $+$ lensing $+$ BAO [94] and could be further probed by upcoming surveys like Euclid+CMB-S4+LiteBIRD [95]. While the projected the effective Majorana mass $m_{ee}$ values fall below current KamLAND-Zen [96], next-generation $0\nu\beta\beta$-decay experiments LEGEND-1000 [97] and nEXO [98] may achieve the required sensitivity for definitive tests. In the further, a very precise measurement of the lepton mixing parameters and neutrino masses might provide a good opportunity for probing this model.
@@
We stress that the numerical values of parameters $m_{a}$, $r$ and $\eta$ remain undetermined by residual symmetry constraints. The phase parameter $\eta$ requires explicit model-building efforts for its determination. Upon specifying $\eta$, all neutrino mixing observables and mass eigenvalues become functions solely of $m_{a}$ and $r$ ultimately constrained through experimental measurements of $\Delta m^{2}_{21}$ and $\Delta m^{2}_{31}$. This framework consequently provides unique predictions for three lepton mixing angles, two CP violation phases and the absolute neutrino mass. Our analysis focuses on the minimal configuration $\eta=\frac{3\pi}{5}$, a value chosen for mathematical simplicity and anticipated feasibility in dynamical model implementations. The subsequent numerical calculations systematically explore correlations between input parameters $m_{a}$, $r$ and output observables. We find that the minimum value of $\chi^{2}$ function is $\chi^{2}_{\text{min}}=5.573$ when the free parameters take the following values
@@
$(G_{l}$,$G_{\text{atm}}$,$G_{\text{sol}})$ $(X^{\text{atm}}_{\bm{\bar{3}_{0}}}$,$X^{\text{sol}}_{\bm{\bar{3}_{1}}})$ $P_{l}$ $(G_{l}$,$G_{\text{atm}}$,$G_{\text{sol}})$ $(X^{\text{atm}}_{\bm{\bar{3}_{0}}}$,$X^{\text{sol}}_{\bm{\bar{3}_{1}}})$ $P_{l}$
@@
$\mathcal{W}_{2}$ ($Z^{abd}_{8},Z^{ad^{2}}_{2},Z^{abc}_{2}$) ($\mathbb{1}_{3},\rho_{\bm{\bar{3}_{1}}}(c^{3})$) $P_{123}$ $\mathcal{W}_{3}$ ($Z^{abd}_{8},Z^{acd^{2}}_{3},Z^{bc^{2}d^{2}}_{2}$) ($\rho_{\bm{\bar{3}_{0}}}(d^{2}),\rho_{\bm{\bar{3}_{1}}}(c^{3}d)$) $P_{231}$
Table 2: The 42 viable breaking patterns, where $\mathcal{T}_{i}$, $\mathcal{F}_{i}$, $\mathcal{S}_{i}$ and $\mathcal{W}_{i}$ denote those patterns in which the residual symmetries in the charged lepton sector correspond to the cyclic groups $Z^{ac}_{3}$, $Z^{c}_{4}$, $Z^{abd^{2}}_{4}$ and $Z^{abd}_{8}$, respectively. $P_{l}$ represents the undetermined permutation matrix of PMNS matrix in Eq. (2.21).
@@
For each of the 5,952 possible residual symmetry combinations, it is necessary to evaluate their compatibility with experimental data by determining whether the corresponding symmetry breaking patterns can reproduce observed results within experimental uncertainties. To this end, we perform a comprehensive numerical analysis for all possible breaking patterns. In our analysis, the $\chi^{2}$ statistic depends on four input parameters which are randomly sampled from the ranges $m_{a}\in[0,1]\text{eV}$, $r\in[0,100]$, $x\in[-20,20]$ and $\eta\in[0,2\pi]$, with the goal of identifying the minimum $\chi^{2}$ values. Furthermore, we concentrate on predicting symmetry breaking patterns that yield the three mixing angles $\theta_{13}$, $\theta_{12}$, $\theta_{23}$, the Dirac CP phase $\delta_{CP}$ and the mass squared differences $\Delta m^{2}_{21}$, $m^{2}_{31}$ within their respective $3\sigma$ ranges of global data analysis summarized in table 1. Through a comprehensive $\chi^{2}$ analysis of all 5,952 potential symmetry breaking configurations, we identify 42 distinct patterns whose theoretical predictions closely align with experimental measurements of neutrino mass parameters and mixing angles [^6] We only kept breaking patterns with $\chi^{2}_{\text{min}}\leq 10$, as we think that those with $\chi^{2}_{\text{min}}>10$ do not exhibit particularly good agreement with experiments.. The 42 viable breaking patterns and the corresponding residual symmetry combinations are summarized in table 2. Specifically, among these 42 breaking patterns: 27 feature the charged lepton residual symmetry $G_{l}$ as $Z^{ac}_{3}$ (labeled $\mathcal{T}_{i}$), 7 exhibit $Z^{c}_{4}$ (labeled $\mathcal{F}_{i}$), 5 demonstrate $Z^{abd^{2}}_{4}$ (labeled $\mathcal{S}_{i}$), and 3 display $Z^{abd}_{8}$ (labeled $\mathcal{W}_{i}$). The best fit values of the input parameters, three mixing angles $\theta_{12}$, $\theta_{13}$, $\theta_{23}$, Dirac CP violation phase $\delta_{CP}$, Majorana CP violation phase $\beta$, the light neutrino masses $m_{2}$, $m_{3}$ and the effective mass $m_{ee}$ in $0\nu\beta\beta$-decay at the minimum values of $\chi^{2}$ function for 42 viable breaking patterns are summarized in table 3. The predicted values of $m_{ee}$ across all viable symmetry breaking patterns fall within the range $[2\,\text{eV},4\,\text{eV}]$, which aligns with the KamLAND-Zen experiment’s latest upper limit of $m_{ee}<(28-122)\,\text{meV}$ [96]. These predictions also lie near the sensitivity thresholds of next generation detectors such as LEGEND-1000 (targeting $m_{ee}<(9\sim 21)\,$meV) [97] and nEXO (projected to reach $m_{ee}<(4.7\sim 20.3)\,$meV) [98]. After scanning the parameter space of $m_{a}$, $r$, $x$ and $\eta$ we identified their viable ranges consistent with neutrino oscillation data at $3\sigma$ level, yielding predictions for lepton mixing parameters and neutrino masses, please see table LABEL:tab:viable_regions.
@@
Furthermore, we can pin down the minimal value of $\chi^{2}$ for each breaking pattern relevant to three lepton mixing angles, and two CP violation phases. These numerical outcomes are graphically summarized in figure 3. Analysis of the figure reveals that almost half of the theoretically permissible breaking patterns yield optimal parameter values for the mixing angles and Dirac CP phase that reside within the experimentally constrained $1\sigma$ confidence intervals reported in Ref. [50]. These breaking patterns can agree with the experimental data very well. Theoretically, advanced measurements of neutrino mass properties and lepton mixing parameters from next-generation neutrino oscillation experiments and cosmological surveys could enable rigorous testing of different breaking patterns. Particularly noteworthy is the JUNO experiment [91] projected to achieve sub-percent precision in determining the solar mixing angle $\theta_{12}$. Figure 3 illustrates the anticipated $3\sigma$ confidence intervals for $\sin^{2}\theta_{12}$ following six years of operational data collection at JUNO, highlighting its potential to constrain theoretical models. The feasibility of these symmetry breaking patterns will undergo rigorous evaluation through upcoming high precision measurements of $\theta_{23}$ and $\delta_{CP}$ parameters conducted by next generation neutrino experiments DUNE [92] and T2HK [93]. Figure 3 illustrates the angular resolution capabilities achieved after 15 years of data acquisition in the DUNE experiment. The collaborative interplay between JUNO and the long baseline neutrino experiments DUNE and T2HK demonstrates remarkable potential for investigating almost all of these breaking patterns.
@@
Figure 3: Best fit results for 42 viable breaking patterns showing $\chi^{2}$ minima, lepton mixing angles, and CP violation phases. Red dashed lines indicate best fit values, light blue bands show the $1\sigma$ and $3\sigma$ ranges from NuFIT 6.0 with Super-Kamiokande atmospheric data [50]. The pale green band shows the projected $3\sigma$ range for $\sin^{2}\theta_{12}$ after 6 years of JUNO data [91]. The faint green areas indicate the resolution (in degrees) for $\sin^{2}\theta_{23}$ and $\delta_{CP}$ after 15 years of DUNE operation [92].
@@
$\mathcal{W}_{3}$ $-0.918$ 0.920 6.26 2.18 4.30 0.02216 0.312 0.466 1.47 1.71 8.66 50.1 2.38
Table 3: The best fit results for input parameters and lepton observables for the 42 viable phenomenologically interesting breaking patterns in table 2.
Table 4: The predicted ranges of the input parameters and lepton observables for all the 42 viable breaking patterns. Note the mixing angle $\theta_{13}$ can take any values within their $3\sigma$ range, and the neutrino mass $m_{3}$ is predicted to lie in $[49.5\text{meV},50.8\text{meV}]$ for each breaking pattern. We would not show them here.
@@
$\mathcal{W}_{3}$ [--1.05,--0.876] [0.763,0.970] [6.01,6.45] [1.61,2.38] [0.303,0.345] [0.435,0.556] [1.44,1.53] [1.47,1.78] [8.32,8.97] [2.17,3.15]

### v2

@@ 3.1 Lepton mixing patterns for $\phi_{\text{atm}},\phi_{\text{sol}}\sim\mathbf{\bar{3}_{0}}$
Moreover, we plot the contour regions for the $3\sigma$ intervals of mixing parameters $\theta_{13}$, $\theta_{23}$, $\delta_{CP}$ and mass ratio $m^{2}_{2}/m^{2}_{3}$ in the plane $r$ and $\eta$ in figure 1. The result for $\theta_{12}$ is not displayed here, because it is related to the reactor angle $\theta_{13}$ by the TM1 mixing sum rule in Eq. (3.12). The JUNO experiment [95], an upcoming medium-baseline reactor neutrino facility, is projected to constrain $\sin^{2}\theta_{12}$ within $[0.3022,0.3118]$ at $3\sigma$ after six years of operation. Complementary tests of $\sin^{2}\theta_{23}$ and $\delta_{CP}$ resolutions (anticipated ranges of $[0.462,0.480]$ and $[1.099,1.255]$ in radians after 15 years, respectively) will be enabled by long-baseline experiments DUNE [96] and T2HK [97]. The predicted neutrino mass sum $\sum_{i=1}^{3}m_{i}$ aligns with cosmological constraints from Planck ++ lensing ++ BAO [98] and could be further probed by upcoming surveys like Euclid+CMB-S4+LiteBIRD [99]. The predicted effective Majorana mass $m_{ee}$ values fall below current KamLAND-Zen [100].Simultaneously, it also lies below the projected sensitivities of next-generation $0\nu\beta\beta$-decay experiments LEGEND-1000 [101] and nEXO [102]. Furthermore, a very precise measurement of the lepton mixing parameters and neutrino masses might provide a good opportunity for probing this model.
@@
We stress that the numerical values of parameters $m_{a}$, $r$ and $\eta$ remain undetermined by residual symmetry constraints. The phase parameter $\eta$ requires explicit model-building efforts for its determination. Upon specifying $\eta$, all neutrino mixing observables and mass eigenvalues become functions solely of $m_{a}$ and $r$ ultimately constrained through experimental measurements of $\Delta m^{2}_{21}$ and $\Delta m^{2}_{31}$. This framework consequently provides unique predictions for three lepton mixing angles, two CP violation phases and the absolute neutrino masses. Our analysis focuses on the minimal configuration $\eta=\frac{3\pi}{5}$, a value chosen for mathematical simplicity and anticipated feasibility in dynamical model implementations. The subsequent numerical calculations systematically explore the correlations between input parameters $m_{a}$, $r$ and output observables. We find that the minimum value of $\chi^{2}$ function is $\chi^{2}_{\text{min}}=5.573$ when the free parameters take the following values
@@
For the NO case
$(G_{l}$$G_{\text{atm}}$$G_{\text{sol}})$ $(X^{\text{atm}}_{\bm{\bar{3}_{0}}}$$X^{\text{sol}}_{\bm{\bar{3}_{1}}})$ $P_{l}$ $(G_{l}$$G_{\text{atm}}$$G_{\text{sol}})$ $(X^{\text{atm}}_{\bm{\bar{3}_{0}}}$$X^{\text{sol}}_{\bm{\bar{3}_{1}}})$ $P_{l}$
@@
$\mathcal{W}_{2}$ ($Z^{abd}_{8},Z^{ad^{2}}_{3},Z^{abc}_{2}$) ($\mathbb{1}_{3},\rho_{\bm{\bar{3}_{1}}}(c^{3})$) $P_{123}$ $\mathcal{W}_{3}$ ($Z^{abd}_{8},Z^{acd^{2}}_{3},Z^{bc^{2}d^{2}}_{2}$) ($\rho_{\bm{\bar{3}_{0}}}(d^{2}),\rho_{\bm{\bar{3}_{1}}}(c^{3}d)$) $P_{231}$
For the IO case
$\mathcal{T}_{1}^{\prime}$ ($Z^{ac}_{3},Z^{c^{2}}_{2},Z^{a^{2}b}_{2}$) ($\mathbb{1}_{3},\rho_{\bm{\bar{3}_{1}}}(c^{2}d^{3})$) $P_{231}$ $\mathcal{T}_{2}^{\prime}$ ($Z^{ac}_{3},Z^{a}_{3},Z^{a^{2}bd^{2}}_{2}$) ($\mathbb{1}_{3},\rho_{\bm{\bar{3}_{1}}}(c^{2}d)$) $P_{321}$
$\mathcal{T}_{3}^{\prime}$ ($Z^{ac}_{3},Z^{ac^{2}}_{3},Z^{a^{2}bd^{3}}_{2}$) ($\mathbb{1}_{3},\rho_{\bm{\bar{3}_{1}}}(d)$) $P_{312}$ $\mathcal{T}_{4}^{\prime}$ ($Z^{ac}_{3},Z^{ac^{2}}_{3},Z^{a^{2}bd^{3}}_{2}$) ($\mathbb{1}_{3},\rho_{\bm{\bar{3}_{1}}}(c^{2}d^{2})$) $P_{321}$
$\mathcal{T}_{5}^{\prime}$ ($Z^{ac}_{3},Z^{acd}_{3},Z^{abc^{3}}_{2}$) ($\rho_{\bm{\bar{3}_{0}}}(bc^{3}d),\rho_{\bm{\bar{3}_{1}}}(d^{2})$) $P_{312}$ $\mathcal{T}_{6}^{\prime}$ ($Z^{ac}_{3},Z^{acd^{2}}_{3},Z^{b}_{2}$) ($\rho_{\bm{\bar{3}_{0}}}(d^{2}),\mathbb{1}_{3}$) $P_{312}$
$\mathcal{F}_{1}^{\prime}$ ($Z^{c}_{4},Z^{a}_{3},Z^{a^{2}bd}_{2}$) ($\mathbb{1}_{3},\rho_{\bm{\bar{3}_{1}}}(c^{2}d^{2})$) $P_{123}$ $\mathcal{F}_{2}^{\prime}$ ($Z^{c}_{4},Z^{ac^{2}}_{3},Z^{abc^{3}}_{2}$) ($\mathbb{1}_{3},\rho_{\bm{\bar{3}_{1}}}(d^{2})$) $P_{321}$
$\mathcal{S}_{1}^{\prime}$ ($Z^{abd^{2}}_{4},Z^{a^{2}bd}_{2},Z^{b}_{2}$) ($\rho_{\bm{\bar{3}_{0}}}(c^{2}d^{2}),\mathbb{1}_{3}$) $P_{321}$ $\mathcal{S}_{2}^{\prime}$ ($Z^{abd^{2}}_{4},Z^{a}_{3},Z^{bc^{2}d^{2}}_{2}$) ($\mathbb{1}_{3},\rho_{\bm{\bar{3}_{1}}}(c^{3}d)$) $P_{231}$
$\mathcal{W}_{1}^{\prime}$ ($Z^{abd}_{8},Z^{a^{2}b}_{2},Z^{b}_{2}$) ($\mathbb{1}_{3},\rho_{\bm{\bar{3}_{1}}}(cd^{3})$) $P_{321}$ $\mathcal{W}_{2}^{\prime}$ ($Z^{abd}_{8},Z^{ac^{2}}_{3},Z^{abc^{3}}_{2}$) ($\mathbb{1}_{3},\rho_{\bm{\bar{3}_{1}}}(c^{3})$) $P_{132}$
Table 2: The 54 viable breaking patterns, where $\mathcal{T}_{i}$ ($\mathcal{T}_{i}^{\prime}$), $\mathcal{F}_{i}$ ($\mathcal{F}_{i}^{\prime}$), $\mathcal{S}_{i}$ ($\mathcal{S}_{i}^{\prime}$) and $\mathcal{W}_{i}$ ($\mathcal{W}_{i}^{\prime}$) denote those patterns in which the residual symmetries in the charged lepton sector correspond to the cyclic groups $Z^{ac}_{3}$, $Z^{c}_{4}$, $Z^{abd^{2}}_{4}$ and $Z^{abd}_{8}$, respectively. $P_{l}$ represents the undetermined permutation matrix of PMNS matrix in Eq. (2.21).
@@
For each of the 5,952 possible residual symmetry combinations, it is necessary to evaluate their compatibility with experimental data by determining whether the corresponding symmetry breaking patterns can reproduce observed results within experimental uncertainties. To this end, we perform a comprehensive numerical analysis for all possible breaking patterns. In our analysis, the $\chi^{2}$ statistic depends on four input parameters which are randomly sampled from the ranges $m_{a}\in[0,1]\text{eV}$, $r\in[0,100]$, $x\in[-20,20]$ and $\eta\in[0,2\pi]$, with the goal of identifying the minimum $\chi^{2}$ values. Furthermore, we concentrate on predicting symmetry breaking patterns that yield the three mixing angles $\theta_{13}$, $\theta_{12}$, $\theta_{23}$, the Dirac CP phase $\delta_{CP}$ and the mass squared differences $\Delta m^{2}_{21}$, $\Delta m^{2}_{31}$ ($\Delta m^{2}_{32}$) for the NO (IO) case within their respective $3\sigma$ ranges of global data analysis summarized in table 1. Through a comprehensive $\chi^{2}$ analysis of all 5,952 potential symmetry breaking configurations, we identify 42 (12) distinct patterns whose theoretical predictions closely align with experimental measurements of neutrino mass parameters and mixing angles in NO (IO) case [^6] We only kept breaking patterns with $\chi^{2}_{\text{min}}\leq 10$, as we think that those with $\chi^{2}_{\text{min}}>10$ do not exhibit particularly good agreement with experiments.. The 54 viable breaking patterns and the corresponding residual symmetry combinations are summarized in table 2. Specifically, among these 54 breaking patterns: 33 feature the charged lepton residual symmetry $G_{l}$ as $Z^{ac}_{3}$ (labeled $\mathcal{T}_{i}$ and $\mathcal{T}_{i}^{\prime}$), 9 exhibit $Z^{c}_{4}$ (labeled $\mathcal{F}_{i}$ and $\mathcal{F}_{i}^{\prime}$), 7 demonstrate $Z^{abd^{2}}_{4}$ (labeled $\mathcal{S}_{i}$ and $\mathcal{S}_{i}^{\prime}$), and 5 display $Z^{abd}_{8}$ (labeled $\mathcal{W}_{i}$ and $\mathcal{W}_{i}^{\prime}$). For the NO (IO) case, the best fit values of the input parameters, three mixing angles $\theta_{12}$, $\theta_{13}$, $\theta_{23}$, Dirac CP violation phase $\delta_{CP}$, Majorana CP violation phase $\beta$, the light neutrino masses $m_{2}$, $m_{3}$ ($m_{1}$, $m_{2}$) and the effective mass $m_{ee}$ in $0\nu\beta\beta$-decay at the minimum values of $\chi^{2}$ function for 54 viable breaking patterns are summarized in table LABEL:tab:viable_best_fit. In the case of the NO spectrum, the predicted values of $m_{ee}$ across all viable symmetry breaking patterns fall within the range $2\,\text{meV},4\,\text{meV}]$ which is too small to be detectable by both current and future $0\nu\beta\beta$-decay experiments. These predictions are significantly below the latest KamLAND-Zen bound of $m_{ee}<(28-122)\,\text{meV}$ [100], and also fall below the projected sensitivity thresholds of next-generation experiments such as LEGEND-1000 (targeting $m_{ee}<(9\sim 21)\,$meV) [101] and nEXO (projected to reach $m_{ee}<(4.7\sim 20.3)\,$meV) [102]. In contrast, for the IO spectrum, future $0\nu\beta\beta$-decay experiments are anticipated to be able to test all viable breaking patterns. After scanning the parameter space of $m_{a}$, $r$, $x$ and $\eta$, we identified their viable ranges consistent with neutrino oscillation data at $3\sigma$ level, yielding predictions for lepton mixing parameters and neutrino masses. Please see table LABEL:tab:viable_regions.
@@
Furthermore, we can pin down the minimal value of $\chi^{2}$ for each breaking pattern relevant to three lepton mixing angles, and two CP violation phases. The numerical outcomes for the NO and IO cases are graphically summarized in figures 3 and  4, respectively. Analysis of the figure reveals that almost half of the theoretically permissible breaking patterns yield optimal parameter values for the mixing angles and Dirac CP phase that reside within the experimentally constrained $1\sigma$ confidence intervals reported in Ref. [50]. These breaking patterns can agree with the experimental data very well. Theoretically, advanced measurements of neutrino mass properties and lepton mixing parameters from next-generation neutrino oscillation experiments and cosmological surveys could enable rigorous testing of different breaking patterns. Particularly noteworthy is the JUNO experiment [95] predicted to achieve sub-percent precision in determining the solar mixing angle $\theta_{12}$. Figures 3 and  4 illustrate the anticipated $3\sigma$ confidence intervals for $\sin^{2}\theta_{12}$ following six years of operational data collection at JUNO, highlighting its potential to constrain theoretical models. The feasibility of these symmetry breaking patterns will undergo rigorous evaluation through upcoming high precision measurements of $\theta_{23}$ and $\delta_{CP}$ parameters conducted by next generation neutrino experiments DUNE [96] and T2HK [97]. Figures 3 and  4 also illustrate the angular resolution capabilities achieved after 15 years of data acquisition in the DUNE experiment. The collaborative interplay between JUNO and the long baseline neutrino experiments DUNE and T2HK demonstrates remarkable potential for investigating almost all of these breaking patterns.
@@
Figure 3: Best fit results for 42 viable breaking patterns showing $\chi^{2}$ minima, lepton mixing angles, and CP violation phases for the NO case. Red dashed lines indicate best fit values, light blue bands show the $1\sigma$ and $3\sigma$ ranges from NuFIT 6.0 with Super-Kamiokande atmospheric data [50]. The pale green band shows the predicted $3\sigma$ range for $\sin^{2}\theta_{12}$ after 6 years of JUNO data [95]. The faint green areas indicate the resolution (in degrees) for $\sin^{2}\theta_{23}$ and $\delta_{CP}$ after 15 years of DUNE operation [96].
Figure 4: The the best fit $\chi^{2}$ for all 12 viable breaking patterns with IO neutrino mass spectrum, covering the three lepton mixing angles and CP violation phases.
Table 3: The best fit results for input parameters and lepton observables for the 54 viable phenomenologically interesting breaking patterns in table 2.
For the NO case
@@
$\mathcal{W}_{3}$ 0.918-0.918 0.920 6.26 2.18 4.30 0.02216 0.312 0.466 1.47 1.71 8.66 50.1 2.38
For the IO case
$x$ $\eta/\pi$ $\frac{m_{a}}{\text{meV}}$ $r$ $\chi^{2}_{\text{min}}$ $\sin^{2}\theta_{13}$ $\sin^{2}\theta_{12}$ $\sin^{2}\theta_{23}$ $\delta_{CP}/\pi$ $\beta/\pi$ $\frac{m_{1}}{\text{meV}}$ $\frac{m_{2}}{\text{meV}}$ $\frac{m_{ee}}{\text{meV}}$
$\mathcal{T}_{1}^{\prime}$ 0.543 1.50 57.6 0.464 3.94 0.02232 0.308 0.574 1.82 1.67 49.1 49.8 42.9
$\mathcal{T}_{2}^{\prime}$ 0.0938 0.00455 20.1 2.87 0.0483 0.02230 0.308 0.553 1.17 1.88 49.1 49.8 47.5
$\mathcal{T}_{3}^{\prime}$ 0.153 0.993 18.0 2.71 4.13 0.02222 0.308 0.574 1.53 0.158 49.1 49.8 47.0
$\mathcal{T}_{4}^{\prime}$ 0.208 1.83 18.4 2.80 0.685 0.02232 0.308 0.538 1.73 0.0642 49.1 49.8 48.0
$\mathcal{T}_{5}^{\prime}$ -1.42 0.75 25.0 0.757 2.05 0.02224 0.308 0.567 1.34 0.244 49.1 49.8 45.2
$\mathcal{T}_{6}^{\prime}$ 7.12 1.89 19.1 0.0564 0.435 0.02232 0.308 0.540 1.50 1.08 49.1 49.8 19.0
$\mathcal{F}_{1}^{\prime}$ -0.165 0.130 18.9 2.82 2.54 0.02225 0.308 0.569 1.56 1.50 49.1 49.8 36.6
$\mathcal{F}_{2}^{\prime}$ 7.09 1.93 21.8 0.0573 6.44 0.02217 0.308 0.580 1.70 1.64 49.1 49.8 41.7
$\mathcal{S}_{1}^{\prime}$ 6.46 1.10 34.4 0.0461 3.28 0.02234 0.308 0.523 1.26 0.511 49.1 49.8 36.0
$\mathcal{S}_{2}^{\prime}$ 0.382 0 18.7 1.87 1.51 0.02234 0.318 0.561 1.47 1.99 49.1 49.8 48.2
$\mathcal{W}_{1}^{\prime}$ 0.122 0.0845 31.0 1.91 5.09 0.02234 0.308 0.516 1.28 0.609 49.1 49.8 31.5
$\mathcal{W}_{2}^{\prime}$ -3.55 0.995 21.8 0.0582 7.12 0.02240 0.308 0.582 1.50 1.96 49.1 49.8 48.1
Table 4: The predicted ranges of the input parameters and lepton observables for all the 54 viable breaking patterns. Note the mixing angle $\theta_{13}$ can take any values within their $3\sigma$ range, and the neutrino mass $m_{3}$ is predicted to lie in $49.5\text{meV},50.8\text{meV}]$ for the NO case. The mixing angle $\theta_{12}$ can take any values within their $3\sigma$ range, and the neutrino mass $m_{1}$ is predicted to lie in $48.4\text{meV},49.8\text{meV}]$ for the IO case. We would not show them here.
For the NO case
@@
$\mathcal{W}_{3}$ [-1.05,-0.876] [0.763,0.970] [6.01,6.45] [1.61,2.38] [0.303,0.345] [0.435,0.556] [1.44,1.53] [1.47,1.78] [8.32,8.97] [2.17,3.15]
For the IO case
$x$ $\eta/\pi$ $m_{a}$/meV $r$ $\sin^{2}\theta_{13}$ $\sin^{2}\theta_{23}$ $\delta_{CP}/\pi$ $\beta/\pi$ $m_{2}$/meV $m_{ee}$/meV
$\mathcal{T}_{1}^{\prime}$ [0.532,0.554] [1.495,1.497] [56.61,58.53] [0.454,0.475] [0.0207,0.0239] [0.574,0.574] [1.793,1.849] [1.658,1.683] [49.2,50.5] [42.3,43.5]
$\mathcal{T}_{2}^{\prime}$ [0.0906,0.0972] [0.00376,0.00539] [19.89,20.34] [2.860,2.879] [0.0209,0.0238] [0.551,0.554] [1.168,1.179] [1.875,1.884] [49.2,50.4] [46.9,48.1]
$\mathcal{T}_{3}^{\prime}$ [0.146,0.160] [0.992,0.994] [17.64,18.24] [2.689,2.736] [0.0207,0.0240] [0.572,0.577] [1.524,1.538] [0.153,0.162] [49.2,50.5] [46.3,47.7]
$\mathcal{T}_{4}^{\prime}$ [0.199,0.216] [1.827,1.842] [18.09,18.68] [2.779,2.819] [0.0207,0.0237] [0.537,0.538] [1.725,1.733] [0.0527,0.0746] [49.2,50.5] [47.3,48.8]
$\mathcal{T}_{5}^{\prime}$ [-1.434,-1.408] [0.749,0.750] [24.65,25.32] [0.747,0.761] [0.0216,0.0240] [0.566,0.570] [1.332,1.351] [0.234,0.253] [49.2,50.5] [44.6,45.8]
$\mathcal{T}_{6}^{\prime}$ [7.060,7.180] [1.890,1.899] [18.79,19.30] [0.0552,0.0574] [0.0220,0.0226] [0.5398,0.5404] [1.165,1.635] [0.834,1.170] [49.2,50.4] [18.5,19.3]
$\mathcal{F}_{1}^{\prime}$ [-0.172,-0.158] [0.123,0.137] [18.59,19.16] [2.801,2.832] [0.0206,0.0240] [0.567,0.572] [1.529,1.587] [1.481,1.523] [49.2,50.5] [36.0,37.2]
$\mathcal{F}_{2}^{\prime}$ [6.960,7.180] [1.926,1.928] [21.46,22.04] [0.0558,0.0593] [0.0216,0.0230] [0.579,0.582] [1.686,1.721] [1.620,1.647] [49.2,50.5] [41.1,42.3]
$\mathcal{S}_{1}^{\prime}$ [6.353,6.544] [1.093,1.097] [33.94,34.88] [0.0450,0.0476] [0.0218,0.0231] [0.522,0.524] [1.231,1.285] [0.495,0.532] [49.2,50.5] [35.5,36.5]
$\mathcal{S}_{2}^{\prime}$ [0.372,0.392] [0,0.000571] [18.40,18.97] [1.831,1.909] [0.0207,0.0240] [0.558,0.564] [1.450,1.487] [1.991,1.993] [49.2,50.5] [47.5,49.0]
$\mathcal{W}_{1}^{\prime}$ [0.117,0.128] [0.0795,0.08934] [30.46,31.56] [1.896,1.919] [0.0207,0.0240] [0.515,0.517] [1.242,1.327] [0.584,0.640] [49.2,50.5] [31.0,32.0]
$\mathcal{W}_{2}^{\prime}$ [-3.600,-3.502] [0.9947,0.9953] [21.50,22.11] [0.0565,0.0597] [0.0219,0.0230] [0.581,0.583] [1.492,1.497] [1.952,1.957] [49.2,50.5] [47.4,48.8]